OpenStudy (anonymous):

Evaluate the integral 1/[x^(2/3)*(9+x^(1/3))] How to solve it?

OpenStudy (anonymous):

Try \(u=9+x^{1/3}\)

OpenStudy (anonymous):

\[ du=\frac 13 x^{-2/3}dx \]

OpenStudy (anonymous):

\[ \int \frac 1{9+x^{1/3}}(x^{-2/3}dx)\to\int \frac 1u\;(3du) \]

OpenStudy (anonymous):

ah ok i got it now, thank you so much ^_________^

OpenStudy (anonymous):

Prodigious!